OpenStudy (anonymous):

The equation V=-3000t+21000 can be used to find the Value V of a car at the end of t years...how much did the car orginally cost?

OpenStudy (anonymous):

t = 0 (at the time of purchase) V = -3000(0) + 21000 V = 21000

OpenStudy (anonymous):

THANK YOU!